OpenStudy (anonymous):

Which properties are used to add these complex numbers? Check all that apply. (4+3i)+(8+2i)=(4+8i)+(3i+2i)=12+5i A. Associative property B. Commutative property C. Distributive property D. None of the above

OpenStudy (cwrw238):

Commutative property is one property you have used . Its (4 + 8) by the way not ( 4 +8i)

Miracrown (miracrown):

You need commutativity to switch the order of the additions

Miracrown (miracrown):

then you used associativity to group the real numbers with themselves and the complex numbers with themselves and then add them

Miracrown (miracrown):

(4+3i)+(8+2i) = (4)+(3i+8)+(2i) by associativity = (4) + (8+3i) + (2i) by commutativity = (4+8)+(3i+2i) by associativity = 12 + 5i by addition

Miracrown (miracrown):

check all that apply A&B

OpenStudy (anonymous):

so a and b?

OpenStudy (cwrw238):

yes